OpenStudy (anonymous):

21/49,56/9 this doesn't add up ive tried:{ product in lowest terms how do i do this?

OpenStudy (bibby):

start by reducing the first fraction then find the lcm

OpenStudy (dan815):

u wanna multiply those 2 fractions?

OpenStudy (unklerhaukus):

The product of 21/49, with 56/9 \[=\frac{21}{49}\times\frac{56}{9}\\=\frac{21\times56}{49\times9}\\ =\\ =\]

OpenStudy (unklerhaukus):

\[=\frac{21\times56}{49\times9}\\ =\frac{3\times7\times7\times8}{7\times7\times3\times3}\\ =\\ =\]

OpenStudy (wolf1728):

UnkleRhaukus, that's just what I was going to type LOL MaeMoore is till online - hope she sees this.

OpenStudy (unklerhaukus):

\[ =\frac{\color{red}{\cancel{\color{black}{3\times7\times7}}}\times8} {\color{red}{\cancel{\color{black}{7\times7\times3}}}\times3}\\ =\\ =\]

OpenStudy (bibby):

oh, I thought she meant add up as in add the 2 fractions lol